# Statistical Learning of Distributionally Robust Stochastic Control in Continuous State Spaces

### v1

@@ 1 Introduction
We consider managing a portfolio of $m$ assets. $X_{t}\in\mathbb{R}^{m}$ denotes the portfolio at time $t$, where $X_{t,i}$ is the dollar value of asset $i$ at the beginning of time $t$. We can buy and sell assets at the beginning of each time period. Let $A_{t}\in\mathbb{R}^{m}$ be our decision variables at $t$, representing the dollar values of the trades. The state dynamics is $X_{t+1}=W_{t}(X_{t}+A_{t})$, where $W_{t}=\mathrm{diag}(R_{t})\in\mathbb{R}^{m\times m}$ is the a diagonal matrix of asset returns $R_{t}\in\mathbb{R}^{m}$. Here, $R_{t,i}$ represents the return of the $i$-th asset from period $t$ to period $t+1$.
Example 2 (Service and manufacturing systems).
We consider a simple service and manufacturing system with make-to-order queues. Let $X_{n}$ denote the waiting time of the $n$-th job, $W_{n}$ the inter-arrival time between the $(n+1)$-th and the $n$-th job, and assume each job requires 1 unit of work. Let $A_{n}$ be the service rate chosen by the system manager for the $n$-th job. Then, the system dynamics can be written as $X_{n+1}=(X_{n}+1/{A_{n}}-W_{n})^{+}.$
However, in practice, the i.i.d. assumptions for the joint distributions of $\{W_{t}\}$ will often be violated. Challenges such as a correlated noise process, along with the presence of confounders and non-stationarities in the environment, can significantly deteriorate policy performance in real deployment environments. These issues serve as the motivation for adopting a distributionally robust stochastic control (DRSC) formulation. The DRSC framework promotes policy robustness by setting up a dynamic game where the controller selects an action at each time $t$, while another entity–the adversary–perturbs the distribution of $W_{t}$. The introduction of the adversary serves as a strategic device that quantifies the worst-case risk associated with model misspecifications. Although this formulation is natural to dynamic robust decision-making, the learning aspect within DRSC, particularly in continuous state space settings, has not yet been studied. To our knowledge, this paper presents the first optimal sample complexity results for learning in such settings. [^1] It is important to note that the field of Distributionally Robust Reinforcement Learning (DRRL), which is closely related, is quite active. While discussing the differences, it is crucial to highlight two key distinctions: firstly, the function $f$ is known in our context, and secondly, much of the DRRL literature focuses on discrete state-action spaces.
Unlike the robust control literature, which often leads to deterministic optimal controls (González-Trejo et al., 2002), our study explores two distinct DRSC formulations where the adversary can be either current-action-aware (CAA) or current-action-unaware (CAU). The key difference between these models lies in whether the adversary has access to the controller’s realized action when deciding how to perturb $W_{t}$. The presence of deterministic optimal Markov controls and a corresponding Bellman equation are not guaranteed under these conditions, especially when asymmetric information structures are present (Wang et al., 2023c). Specifically, DRSC problems with CAU adversaries necessitate randomized policies to achieve optimal control. While these issues have been explored within the distributionally robust MDP (DRMDP) context (Wiesemann et al., 2013; Wang et al., 2023c), there has been limited research focusing on differentiating between CAA and CAU formulations in robust control settings, or on their implications for learning. Nevertheless, recognizing these distinctions is crucial from modeling, learning, and optimization perspectives. CAU formulations typically result in less powerful adversaries, leading to less conservative controls, although they may be challenging to learn and optimize. Conversely, the more conservative CAA formulations are better suited for highly competitive environments. These modeling considerations are exemplified and discussed in Section 5.
Our formulation contribution extends to establishing the existence of dynamic programming principles, expressed as DR Bellman equations, for both CAA and CAU adversaries, which guarantee the optimality of stationary Markov policies in the discounted infinite horizon setting. Although this contribution is fundamental, it is presented in the supplemental materials due to space constraints. Establishing this dynamic programming principle lays the groundwork for the other main focus of our paper: the statistical complexity of learning the DRSC values, which equal the solution of the corresponding Bellman equations, across a continuous state space.
Our results on learning complexities involve two types of admissible adversarial decisions affecting the distribution of $\{W_{t}:t\geq 0\}$, characterized by Wasserstein distance and $f_{k}$-divergence ambiguity sets. These represent the main types of distributional ambiguity in the field of distributionally robust optimization. Wasserstein ambiguity sets effectively capture model errors at the outcome level, while $f_{k}$-divergence sets hedge against misspecifications in the likelihood of possible outcomes. In Table 1, we provide a summary of the sample complexity results for each of the four cases studied in the paper. The cases are based on the visibility of the current action to the adversary (CAA vs. CAU) and the type of ambiguity sets (Wasserstein distance vs. $f_{k}$-divergence).
Table 1: Summary of our results on the statistical complexity of achieving minimax learning of the DRSC value function in the uniform norm.
Ambiguity Set Type Action Space Sample Complexity Upper Bound Lower Bound
Wasserstein CAA Continuum $\Theta\left(n^{-1/2}\right)$ Theorem 1 Theorem 5
CAU Finite Theorem 3
$f_{k}$-divergence CAA Continuum $\widetilde{\Theta}\left(n^{-\frac{1}{k^{\prime}\vee 2}}\right)$ [^2] The $\widetilde{\Theta}$ suppress a gap of $\sqrt{\log n}$. Theorem 2 Theorem 6
CAU Finite Theorem 4
1.1 Literature Review
Distributionally robust stochastic control is not a new concept in the literature. Yang (2021) investigate a setting aligned with our current-action-aware formulation, employing a Wasserstein uncertainty set. In contrast, Petersen et al. (2000) consider an uncertainty set based on Kullback–Leibler divergence. For linear systems, where $f$ is linear on $X_{t}$,$A_{t}$, and $W_{t}$, distributionally robust stochastic control has been explored by several authors (Taskesen et al., 2024; Kim and Yang, 2023; Han, 2023; Kotsalis et al., 2021). Nonetheless, existing research predominantly focuses on characterizations of the optimal policies or the development of tractable optimization methods. In our study, we address the statistical complexity associated with the learning problem.
Our work is also closely related to the literature on DRMDPs and distributionally robust reinforcement learning (DRRL). Various formulations are explored in Iyengar (2005); Nilim and El Ghaoui (2005); González-Trejo et al. (2002); Le Tallec (2007); Xu and Mannor (2010); Wiesemann et al. (2013); Wang et al. (2023c); Goyal and Grand-Clément (2023); Li and Shapiro (2023). Statistical complexities for associated DRRL problems are subsequently developed in Zhou et al. (2021); Panaganti and Kalathil (2021); Yang et al. (2021); Shi and Chi (2022); Xu et al. (2023); Shi et al. (2023); Blanchet et al. (2023); Liu et al. (2022); Wang et al. (2023a, b); Yang et al. (2023). However, our formulation of DRSC differs as we assume a known state recursion form $f$ driven by an unknown random variable $W$, whereas in DRMDP, the full transition probabilities need to be learned. This difference leads to a significant difference in sample complexities. We can derive minimax optimal sample complexities, even in continuous state and action spaces.
Remarks on Paper Organizations
The remainder of this paper is structured as follows: Section 2 outlines the formulations of the DRSC problems and the corresponding dynamic programming principles, with a focus on both CAA and CAU adversaries. Sections 3 and 4 present the upper and lower bounds of the sample complexities, respectively. Finally, Section 5 presents two applications within our framework to demonstrate its effectiveness in modeling real-world problems.
2 Dynamic Programming and Bellman Equation
In this section, we provide a minimal self-consistent introduction to our formulation of the distributionally robust stochastic optimal control problem and its corresponding dynamic programming theory. The fully rigorous construction is provided in Appendix A. Our focus here is on the infinite horizon discounted reward setting. It should be noted that a DRSC formulation for finite horizon systems naturally arises from the same principles.
We consider Polish (i.e. complete separable metric spaces) state, action, and noise measurable spaces $(\mathbb{X},\mathcal{X}),(\mathbb{A},\mathcal{A}),(\mathbb{W},\mathcal{W})$ equipped with the Borel $\sigma$-fields generated by open sets. Let $\mathcal{P}(\mathcal{W})$ and $\mathcal{P}(\mathcal{A})$ denote the set of probability measures on the action and noise spaces, endowed with the topology of weak convergence. We also consider a known state dynamic function $f:\mathbb{X}\times\mathbb{A}\times\mathbb{W}\rightarrow\mathbb{X}$ given in (1.1), a known reward function $r:\mathbb{X}\times\mathbb{A}\rightarrow\mathbb{R}_{+}$, and a discount factor $\alpha\in(0,1)$.
Let $\mathcal{Q}\subset\mathcal{P}(\mathcal{A})$ and $\mathcal{P}\subset\mathcal{P}(\mathcal{W})$ be arbitrary Borel measurable subsets. Here, $\mathcal{Q}$ and $\mathcal{P}$ represent the admissible decision sets of the controller and the adversary, respectively. Based on $\mathcal{Q}$ and $\mathcal{P}$, we construct admissible policy classes $\Pi(\mathcal{Q})$ and $\Gamma(\mathcal{P})$ of the controller and adversary, respectively. As we will rigorously develop in the Appendix A, the admissible control policy class $\Pi(\mathcal{Q})$ of the controller will always be a proper subset of the history-dependent $\mathcal{Q}$-constrained policy class:
Intuitively, the controller decides a sequence of the conditional distribution of current action $A_{t}$ given the history until the last visible state $x_{t}$.
$\Gamma(\mathcal{P})\subset\Gamma_{\mathrm{H}}:=\left\{{\gamma=(\gamma_{0},\gamma_{1},\cdots):\gamma_{t}(dw|g_{t})\in\mathcal{P},\forall g_{t}=(x_{0},a_{0},\cdots,x_{t},a_{t})}\right\}.$
The adversarial policy $\pi\in\Gamma(\mathcal{P})$ determines the conditional distribution of $W_{t}$ given the history until the last visible state action pair $x_{t},a_{t}$, for each and every $t\geq 0$.
For any given pair of controller and adversarial policy $(\pi,\gamma)\in\Pi(\mathcal{Q})\times\Gamma(\mathcal{P})$ and an initial distribution $\mu\in\mathcal{P}(\mathcal{X})$, the distribution of the state and action process $(X,A)$ is uniquely determined, see (A.2). We denote the expectation under this distribution as $E_{\mu}^{\pi,\gamma}$. Then, we define the DRSC value under initial distribution $\mu$ and controller’s and adversarial policy classes $\Pi(\mathcal{Q})$ and $\Gamma(\mathcal{P})$ as
$v^{*}(\mu,\Pi(\mathcal{Q}),\Gamma(\mathcal{P})):=\sup_{\pi\in\Pi(\mathcal{Q})}\inf_{\gamma\in\Gamma(\mathcal{P})}E_{\mu}^{\pi,\gamma}\sum_{t=1}^{\infty}\alpha^{t}r(X_{t},A_{t}).$ (2.1)
In order for the DRSC value to satisfy a dynamic programming principle, i.e. a Bellman equation, regularity conditions for the reward $r$ and the state transition function $f$ are necessary.
Assumption 1.
Assume that $r:\mathbb{X}\times\mathbb{A}\rightarrow\mathbb{R}_{+}$ is non-negative $r_{\vee}$-bounded (i.e. $0\leq r(x,a)\leq r_{\vee},\forall x,a$) uniformly continuous and $f:\mathbb{X}\times\mathbb{A}\times\mathbb{W}\rightarrow\mathbb{X}$ is uniformly continuous. Furthermore, the controller can take on deterministic decisions; i.e. $\mathcal{Q}\supset\left\{{\delta_{\left\{{a}\right\}}:a\in\mathbb{A}}\right\}$.
Remark.
It is straightforward to generalize all the subsequent results in this paper to the case where the reward at time $t$ is $r(X_{t},A_{t},W_{t})$ which depends on the adversarial noise $W_{t}$. Moreover, for the dynamic programming part, the boundedness of reward and uniform continuities can be significantly weakened to the satisfaction of some growth conditions and semi-continuities (González-Trejo et al., 2002). However, to study the minimax statistical complexity, which is one of the main goals of the paper, we need even stronger conditions (e.g. uniform Lipschitz continuity) as in Assumption 2 and 3. Although relaxing the assumptions to be as general as possible while retaining a dynamic programming theory is of great theoretical value, our focus here is to develop a rigorous theoretical framework that allows us to study the statistical complexities of learning a DRSC. Hence, we adopt these more restrictive assumptions that are easy to work with.
Current-Action-Aware Adversary
Observe that, by our construction, a general adversarial decision $\gamma_{t}$ at time $t$ can be dependent on the entire history $g_{t}=(x_{0},a_{0},\cdots,x_{t},a_{t})$. In particular, the conditional distribution of $W_{t}$ given history and different realizations of $A_{t}$ could be different. In other words, the adversary is aware of the current action, hence the name current-action-aware adversary, and can use this information to harm the performance of the controller. This could be a reasonable model for settings in which ambiguity or non-Markov response arises from the system’s reaction to the control inputs. One such application setting of particular practical importance is the portfolio optimization problem, where policy robustness is extremely valuable.
Definition 1 (Current-Action-Aware Bellman Equation).
We define the following fixed point equation as the Bellman equation for current-action-aware adversaries.
$u(x)$ $=\sup_{\phi\in\mathcal{Q}}\int_{\mathbb{A}}r(x,a)+\alpha\inf_{\psi\in\mathcal{P}}\int_{\mathbb{W}}u(f(x,a,w))\psi(dw)\phi(da)$ (2.2)
$=\sup_{a\in\mathbb{A}}r(x,a)+\alpha\inf_{\psi\in\mathcal{P}}\int_{\mathbb{W}}u(f(x,a,w))\psi(dw)=:\mathcal{T}(u)(x)$
We will show in Proposition 2.1 that, under Assumption 1, the CAA Bellman equation (2.2) has a unique bounded solution $u^{*}$. Moreover, in the full discussion of the DRSC framework in Appendix A, we will rigorously establish the dynamic programming principles under which the DRSC value in (2.1) is equal to $u^{*}$.
Current-Action-Unaware Adversary
Although CAA adversaries give rise to natural distributionally robust control models in various important applications, there are settings for which the noise is considered as exogenous inputs to the system that are independent of the current action of the controller. For example, in the make-to-order manufacturing setting, to achieve robust planning, it might be reasonable to assume the presence of adversarial inter-arrival time given the previous completion times. However, assuming that the inter-arrival times can adversely relate to the service rate could lead to an overly conservative optimal policy, leading to a diminished value in the deployment environment where the interarrival time is independent of the service assignment.
To address this potential issue and increase the versatility of our framework we consider an adversary that cannot observe the current action. Concretely, an adversarial decision $\bar{\gamma}_{t}$ at time $t$ is said to be current-action-unaware (CAU) if for any history $h_{t}=(x_{0},a_{0},\cdots,x_{t})$ and $(h_{t},a)=(x_{0},a_{0},\cdots,x_{t},a)$, we have that $\bar{\gamma}_{t}(dw|h_{t},a)=\bar{\gamma}(dw|h_{t},a^{\prime})$ for all $a,a^{\prime}\in\mathbb{A}$. In other words, the conditional distribution of $W_{t}$ given the history is independent of the current action $A_{t}$. Note that will use the ”overline” notation to signify the CAU setting. We remark that in the DRMDP literature, this behavior is captured by S-rectangular adversaries, see Wiesemann et al. (2013); Wang et al. (2023c) and the reference therein.
For CAU adversaries, we define the following distributionally robust Bellman equation.
Definition 2 (Current-Action-Unaware Bellman Equation).
We define the following fixed point equation as the Bellman equation for current-action-unaware adversaries.
$\bar{u}(x)=\sup_{\phi\in\mathcal{Q}}\inf_{\psi\in\mathcal{P}}\int_{\mathbb{A}\times\mathbb{W}}r(x,a)+\alpha\bar{u}(f(x,a,w))\phi\times\psi(da,dw)=:\overline{\mathcal{T}}(\bar{u})(x)$ (2.3)
Again, we will rigorously establish the dynamic programming principles in the full discussion of the DRSC framework in Appendix A.
[…]
Proposition 2.1.
Suppose Assumption 1 is in force. The Bellman equations (2.2) and (2.3) have unique fixed points $u^{*}$ and $\bar{u}^{*}$ in $U_{b}(\mathbb{X})$, respectively. Moreover, $\left\|u^{*}\right\|,\left\|\bar{u}^{*}\right\|\leq\beta r_{\vee}.$
Combining Proposition 2.1 with the dynamic programming theory in Appendix A.3, we see that learning optimal DRSC value and hence a robust control can be achieved by finding a good approximation to the solutions of (2.2) and (2.3) in CAA and CAU settings, respectively.
3 Upper Bounds on Statistical Complexity
@@
In Definitions 3 and 4, we first introduce Wasserstein distance and $f_{k}$-divergence ambiguity sets. We fix a measure $\mu_{0}\in\mathcal{P}(\mathcal{W})$ as the center of the ambiguity sets, which is only accessible from samples.
@@
For both the Wasserstein distance and $f_{k}$-divergence settings, we let $\widehat{\mathcal{P}}$ do denote the $n$-sample empirical measure-centered version of the ambiguity sets; i.e. one that replaces $\mu_{0}$ with $\hat{\mu}$ where $\hat{\mu}(\cdot):=\frac{1}{n}\sum_{i=1}^{n}\mathds{1}\left\{{W_{i}\in\cdot}\right\}$ for i.i.d. $W_{i}\sim\mu_{0}$. Further, we denote the CAA/CAU empirical Bellman operator as $\mathbf{T}$/$\overline{\mathbf{T}}$, where $\mathcal{P}$ in the definition (2.2)/(2.3) is replaced by $\widehat{\mathcal{P}}$, i.e.,
$\mathbf{T}(u)(x):=\sup_{a\in\mathbb{A}}r(x,a)+\alpha\inf_{\psi\in\widehat{\mathcal{P}}}\int_{\mathbb{W}}u(f(x,a,w))\psi(dw),$
$\overline{\mathbf{T}}(\bar{u})(x):=\sup_{\phi\in\mathcal{Q}}\inf_{\psi\in\widehat{\mathcal{P}}}\int_{\mathbb{A}\times\mathbb{W}}r(x,a)+\alpha\bar{u}(f(x,a,w))\phi\times\psi(da,dw).$
@@
Let $u^{\prime}$ and $\hat{u}$ be the solution to the population and empirical versions of (2.2) or (2.3); let $\mathcal{T}^{\prime}$ and $\mathbf{T}^{\prime}$ denote the corresponding population and empirical Bellman operators. Then, the estimation error in uniform norm is upper bounded by
@@
In the following two sections, we establish statistical complexity upper bounds for learning the optimal DRSC value uniformly under CAA and CAU adversary models. In both cases, we consider both the Wasserstein distance and $f_{k}$-divergence-based adversarial ambiguity set. Our focus here is to obtain a tight convergence rate in $n$.

### v2

@@ 3.3 Modeling Examples
We manage $m$ assets. Let $X_{t}\in\mathbb{R}^{m}$ be the vector of dollar holdings at the start of period $t$ ($X_{t,i}>0$ and $X_{t,i}<0$ mean a long position and a short position in asset $i$, respectively) and let $A_{t}\in\mathbb{R}^{m}$ be resulting percentage of the $m$ assets after trading at $t$ ($A_{t,i}>1$ or $0\leq A_{t,i}<1$ means buying or selling asset $i$ at the beginning of time period $t$, respectively). Let $W_{t}=\mathrm{diag}(R_{t})\in\mathbb{R}^{m\times m}$ be the diagonal matrix of gross returns from $t$ to t+1t{+}1.
For simplicity, we assume that there is no contribution of capital. However, one can consume the portfolio by having a total sale that is higher than the total purchase; i.e. $(A_{t}-\mathbf{1})^{\top}X_{t}\leq 0$ where $\mathbf{1}\in\mathbb{R}^{m}$ is the column vector of all 1’s. Therefore, the consumption is $C_{t}=\mathbf{1}^{\top}X_{t}-A_{t}^{\top}X_{t}\geq 0$, which is the cash amount taken out from the portfolio. As such, the self-financing constraint (in the absence of any transaction costs) becomes $(A_{t}-\mathbf{1})^{\top}X_{t}+C_{t}=0$.
At the beginning of the next period, we have
$X_{t+1}=W_{t}(A_{t}\odot X_{t}).$
We assume that the decision maker is endowed with a utility function $u:\mathbb{R}_{+}\rightarrow\mathbb{R}$, which is concave and non-decreasing. Then, taking action $A_{t}$ will incur an instantaneous reward utility
$r(X_{t},A_{t},W_{t})=u(\mathbf{1}^{\top}X_{t}-A_{t}^{\top}X_{t}).$
CAA vs. CAU: If trades move prices through market impact, the return law effectively reacts to the current trade. Model this with CAA so the adversary can condition $W_{t}$ on $A_{t}$ (temporary impact, slippage, liquidity dry‑ups). If the trader is small relative to the market or executes passively in deep venues, it is reasonable to treat returns as exogenous to the current trade. Then CAU captures serial dependence and regime change without over‑penalizing actions. In short: use CAA to reflect immediate action‑driven risk; use CAU when returns are driven by external factors and impact is negligible.
4 Learning Optimal DRSC: Statistical Complexity Upper Bounds
Beyond modeling expressiveness, a key requirement of the DRSC framework is statistical efficiency: given a reasonably large dataset, we should be able to infer near-optimal policies with reliable fidelity. To provide useful statistical insights and sharp convergence guarantees, it is necessary to tailor the complexity analysis to the specific form of the ambiguity set, since allowing completely general ambiguity sets renders the problem statistically intractable in a minimax sense. In this paper, we therefore focus on two widely used ambiguity models—Wasserstein distance and $f_{k}$-divergence—which capture distinct facets of distributional uncertainty.
When deploying DRSC models, different applications motivate different notions of robustness. At a high level, the design of ambiguity sets for the environment input depends on which aspect of misspecification is most concerning. In some cases, the focus is on outcomes: if realized states or environmental inputs may be perturbed within a small range—due to measurement noise, regime shifts, or model misspecification in the dynamics—it is natural to adopt Wasserstein-type ambiguity sets, which capture closeness in the geometry of the input space [27]. In other cases, the concern lies in likelihoods: the support of outcomes may be trusted, but their relative probabilities may be misspecified, for instance due to sampling bias or confounding. Here, likelihood-based sets, such as those defined by $f_{k}$-divergences, are more suitable because they measure robustness in terms of reweighting likelihood ratios [2]. These two modeling perspectives—outcome perturbations versus likelihood perturbations—motivate the use of Wasserstein and $f_{k}$-divergence ambiguity sets in our framework.
At the same time, Wasserstein or $f_{k}$ ambiguity models enjoy favorable statistical or computational properties. From the distributionally robust optimization literature, Wasserstein balls often yield strong generalization guarantees [12], while $f_{k}$-divergence balls lead to closed-form dual formulations and efficient optimization [8].
@@ 4 Learning Optimal DRSC: Statistical Complexity Upper Bounds
We now formally introduce Wasserstein distance and $f_{k}$-divergence ambiguity sets (Definitions 3 and 4). In both cases, we fix a reference measure $\mu_{0}\in\mathcal{P}(\mathcal{W})$ as the center of the ambiguity set, assumed to be accessible from sampling.
@@
4.1 Data, Learning Objective, and Estimator
To achieve robust and data-driven decision-making, we seek to learn system characteristics and robust decisions from data whose distributional properties may not perfectly align with those in the deployment environment. We assume access to a data set that may be collected either from a simulator of the real system of interest or directly from operations during a data collection period. For analytical tractability, we model this data set as $n$ i.i.d. samples $\{W_{i}:i=1,\ldots,n\}$ with $W_{i}\sim\mu_{0}$.
Our learning objective is to estimate the optimal robust value defined in (3.2) and (3.3) across the entire state space $\mathbb{X}$. Specifically, we aim to construct an estimator $\hat{u}$ of the optimal robust value functions (here denoted by $u$) such that the uniform error
$\|u-\hat{u}\|=\sup_{x\in\mathbb{X}}|u(x)-\hat{u}(x)|$
remains small. This objective requires learning an infinite-dimensional object while ensuring uniform accuracy across the entire state space.
To address this challenge, we adopt the empirical robust Bellman estimator (equivalently, the plug-in estimator in this context), constructed using the empirical measure–centered version of the ambiguity sets introduced below.
For both the Wasserstein distance and $f_{k}$-divergence settings, we let $\widehat{\mathcal{P}}$ do denote the $n$-sample empirical measure-centered version of the ambiguity sets; i.e. one that replaces $\mu_{0}$ with $\hat{\mu}$ where
$\hat{\mu}(\cdot):=\frac{1}{n}\sum_{i=1}^{n}\mathds{1}\left\{{W_{i}\in\cdot}\right\}$
for i.i.d. $W_{i}\sim\mu_{0}$. Further, we denote the CAA/CAU empirical Bellman operator as $\mathbf{T}$$\overline{\mathbf{T}}$, where $\mathcal{P}$ in equation (3.2)/(3.3) is replaced by $\widehat{\mathcal{P}}$, i.e.,
$\mathbf{T}(u)(x):=\sup_{a\in\mathbb{A}}\inf_{\psi\in\widehat{\mathcal{P}}}\int_{\mathbb{W}}r(x,a,w)+\alpha u(f(x,a,w))\psi(dw),$
and
$\overline{\mathbf{T}}(\bar{u})(x):=\sup_{\phi\in\mathcal{Q}}\inf_{\psi\in\widehat{\mathcal{P}}}\int_{\mathbb{A}\times\mathbb{W}}r(x,a,w)+\alpha\bar{u}(f(x,a,w))\phi\times\psi(da,dw).$
@@ 4.1 Data, Learning Objective, and Estimator
Suppose Assumption 1 is in force. Let $u^{\prime}$ and $\hat{u}$ be the solution to the population and empirical versions of (3.2) or (3.3); let $\mathcal{T}^{\prime}$ and $\mathbf{T}^{\prime}$ denote the corresponding population and empirical Bellman operators. Then, the estimation error in uniform norm is upper bounded by
@@
In the following two sections, we establish statistical complexity upper bounds for learning the optimal value uniformly under CAA and CAU adversary models. In both cases, we consider both the Wasserstein distance and $f_{k}$-divergence-based adversarial ambiguity set. Our focus here is to obtain a tight convergence rate in $n$.